# RATSF: Empowering Customer Service Volume Management through Retrieval-Augmented Time-Series Forecasting

### v1

@@
Abstract
An efficient customer service management system hinges on precise forecasting of service volume. In this scenario, where data non-stationarity is pronounced, successful forecasting heavily relies on identifying and leveraging similar historical data rather than merely summarizing periodic patterns. Existing models based on RNN or Transformer architectures often struggle with this flexible and effective utilization. To address this challenge, we propose an efficient and adaptable cross-attention module termed RACA, which effectively leverages historical segments in forecasting task, and we devised a precise representation scheme for querying historical sequences, coupled with the design of a knowledge repository. These critical components collectively form our Retrieval-Augmented Temporal Sequence Forecasting framework (RATSF). RATSF not only significantly enhances performance in the context of Fliggy hotel service volume forecasting but, more crucially, can be seamlessly integrated into other Transformer-based time-series forecasting models across various application scenarios. Extensive experimentation has validated the effectiveness and generalizability of this system design across multiple diverse contexts.
1 Introduction
At the heart of customer service volume management lies the accurate estimation of total service demand, which significantly impacts system costs. For every underestimation of 100 service requests, there is a need to urgently mobilize and supplement service staff at a cost equivalent to three times the single-unit labor cost; conversely, overestimating by 100 requests results in a waste of that same single-unit labor cost. Service volume forecasting for the Travel industry is particularly intricate because of its inherent interconnections with numerous variables, including order statuses, weather variations, international developments, and destination country policies. These factors contribute to highly non-stationary data patterns that, in turn, can lead to considerable biases when employing traditional uni-variate time series forecasting methods [9, 1, 23] which rely on period summaries and trend analyses.
@@ 1 Introduction
One naive way to solve this is to try elongating the sequence that the transformer processes. Some approach involves sampling historical data to fit within a limited context window. Specifically, the Informer [32] algorithm samples K points from the sequence and derives a shorter Q sequence based on these sampled points. However, this approach assumes that all historical information is equally important, which may not be suitable for many time series scenarios where different data points can carry varying significance. Perceiver [10] and similar methods opt for a different approach by mapping Query sequences to fixed lengths, reducing computation and allowing for more historical data storage. Nonetheless, they still face challenges in efficiently extracting and interpreting critical information from extended time series.
In the field of natural language processing, strides have been made to expand Transformer models’ context handling. On one front, techniques like flash-attention [5] enhance efficiency and reduce complexity, enabling longer context processing. However, it’s crucial to note that directly extending receptive fields may cause larger models to overlook significant details in lengthy inputs, like [15] mentioned.
On another front, Retrieval-Augmented Generation (RAG)-like methods have drawn attention for enhancing model performance by incorporating external information. NVIDIA’s [2] research indicates that even when models already handle large context windows in text tasks, they can still achieve substantial performance gains by retrieving and using relevant data from external sources.
Employing the concept of RAG, we have identified a potential approach. We can build a knowledge base that encompasses all historical time series, and pinpoint and utilize the most predictive historical segments with an improved transformer model, thereby refining forecast accuracy. To accomplish these, we focused on implementing three central enhancements: a knowledge base schema for efficiently indexing historical data, a cross-attention module to integrate historical information, and a predictive embedding for retrieval purposes.
Figure 1: RATSF uses a TSKB to store and index historical sequences, alongside a transformer-based forecasting model. As illustrated in the bottom-right corner, TSKB segments the full history for efficient retrieval. To do forecasting, the Encoder fetches $\mathbf{X}_{o}$ from t recent time steps, and forms a retrieval sequence with d latest steps, retrieves N related sequences $\mathbf{X}_{r}$ from the TSKB. RACA in the Decoder then merges $\mathbf{X}_{r}$ and $\mathbf{X}_{o}$ info to deliver result.
@@
1. We offer a general retrieval-augmented time-series forecasting strategy, RATSF, that can be implemented across a variety of Transformer models and their derivatives to enhance performance.
2. We employ a novel embedding strategy for precise historical piece retrieving.
@@
3. We present a straightforward and manageable design for a time-series knowledge base, which significantly facilitates efficient management of historical data.
4. The method presented in this work currently serves as the operational algorithm for Fliggy’s service volume management system across all business domains, supporting not only hotel online bookings but also support businesses such as visa processing, flight and train ticket reservations, as well as entrance tickets to resort areas. Our method has demonstrated exceptional effectiveness in various practical application scenarios.
2 Review
@@
Despite intense competition [20, 6, 30, 8], Transformer models [25] and their improved variants [19, 26, 16, 13, 14] have become the mainstream choice in time-series forecasting tasks. However, the computational complexity of the original Transformer model scales as O(n²) with respect to sequence length, significantly limiting the maximum sequence length it can handle and thereby constraining its ability to use historical data.
@@
Improvements in Time-Series Forecasting with Transformers. Numbers of works focus on enhancing the Transformer model’s capability to extract temporal features, thereby improving prediction accuracy, exemplified by Fedformer’s [33] introduction of a frequency-augmented Attention mechanism that directly incorporates Fourier operators, a similar path taken by TimesNet [27]. In contrast, Autoformer [28] proposes an operator that decomposes time series information into trend and periodic components. These two types of solutions perform well for relatively stationary sequences but may see performance drop when dealing with highly non-stationary time series data.
@@
Another core strategy involves learning a robust representation of the time series first, which is then used to enhance the forecasting accuracy. TNC [24] harnesses the core concept of contrastive learning and employs samples within a specific temporal neighborhood within a window as positive pairs, while treating samples from differing temporal neighborhoods as negative pairs. Despite these improvements having collectively boosted the ability of Transformers in time-series forecasting tasks, none has fundamentally increased the Transformer’s receptive field.
@@
Retrieve Augmented Method in NLP. In last two years, the Retrieval-Augmented Generation (RAG) approach [7, 4, 18, 31, 21, 12, 34] has gained widespread adoption in the field of NLP. REALM harnesses a knowledge retriever to distill information from vast corpora and thereby enhance the performance of pre-trained language models. Transformer-XL+kNN [3] incorporates a K-Nearest Neighbors algorithm to search through training data, refining dialogue generation capabilities. In the context of named entity recognition tasks, U-RaNER [22] utilizes multi-modal heterogeneous retrieval techniques to boost knowledge retrive and, by integrating retrieved knowledge into the model, strengthens its understanding of queries and improves entity recognition accuracy.
@@
Retrieval Augmented Method in Time-series Forecasting. MQ-ReTCNN [29] is designed for complex time series prediction tasks involving multiple entities and variables. It employs a scoring function to compile relevant contexts from offline data, selects scored segments, and appends them to the prediction sequence.Its retrieval mechanism emphasizes leveraging historical sequences of one entity to enhance predictions about another, without directly utilizing historical data for direct prediction assistance.
@@
ReTime [11] creates a relation graph based on temporal closeness between sequences and employs relational retrieval instead of content-based retrieval.It does not optimally use historically similar sequences as reference points due to its inherent design limitations. Both MQ-ReTCNN and ReTime incorporate retrieval enhancement strategies but have yet to introduce a general and efficient retrieval technique specifically for single-variable time series prediction scenarios.
@@
Figure 2: Each RACA has two cross-attention modules: one utilizes the Encoder output as K,V, while the other employs embedded retrieved sequences as K,V. Both outputs are concatenated and passed through a Linear module to reshape back to the input dimensions.
3 Method
Subsequently, we first provide an overview of the RATSF’s overall structure and then examine the key features of each component in detail.
The simplest description of RATSF is to first build a TSKB and then retrieve relevant sequences from it to enhance the forecasting model. For a Time Series Forecasting (TSF) task within a specific context, we first construct a TSKB using the schema detailed in Section 3.2. Next, we utilize a retrieval method to gather relevant sequences, with this method being introduced in Section 3.3. Finally, we integrate these retrieved sequences into the forecasting model, with its core computational component, RACA, being introduced in Section 3.1. Figure 1 shows the data flow of RATSF.
3.1 RACA
@@
As previously mentioned, Retrieval-Augmented Cross-Attention (RACA) is a module designed to be plugged-into a time-series forecasting model. In our demonstration, we employ the vanilla Transformer as the time-series forecasting model; however, it’s important to note that the RACA module is designed to be compatible and can be seamlessly plug into any transformer-based time-series model.
we now provide an explanation of the abbreviations represented by subscript and superscript symbols in the following paragraphs: The letter ’o’ for original, represents the input time-series piece entering the encoder; letter ’b’ for embedding , letter ’e’ for encoder; letter ’r’ for retrieved sequence, letter ’f’ for sequence for forecasting; ’l’ for layer ; ’D’ for hidden dimension.
In transformer architecture, an encoder consists of an embedding module and multiple self-attention modules. Initially, the raw time series data $\mathbf{X}_{o}=\left\{x_{1},...,x_{t}\right\}$, $\mathbf{X}_{o}$ when processed through the embedding module, generates an output as the $\mathbf{X}^{b}_{o}$; subsequently, the output resulting from all the encoder calculation is referred as the $\mathbf{X}^{e}_{o}$.
RACA is introduced within the decoder part: besides conventionally utilizing the encoder output and the place holding embedding for the forecasting sequence, it incorporates information from N most relevant history sequences retrieved from TSKB. Like shown in Figure 2, these retrieved sequences are transformed through the encoder’s embedding module, generating Ns $\mathbf{X}^{e}_{r}$  [^1] In section 4, when comparing four Transformer models horizontally, the impact of the Encoder extends to both representation retrieval and decoding stages. To isolate the contribution of the Encoder to the retrieval performance, we employ the Encoder output as the retrieval scheme while incorporating the representation $\mathbf{X}^{b}_{r}$ rather than $\mathbf{X}^{e}_{r}$ into RACA. In section 3.3, we directly compare the performance of using DTW alone versus Encoder output for retrieval. When integrated into the Decoder, the Encoder output’s affecting both retrieval effectiveness and decoding quality. Given the significant differences between Encoders trained solely with DTW and those following our proposed method,as shown in B.2. To ensure consistency across experiments, $\mathbf{X}^{e}_{r}$ is consistently replaced with $\mathbf{X}^{b}_{r}$ in this paper’s all experiments., and we concatenate these N vectors into one with shape$[1,N*L_{r},D]$ for later use.
@@
In Transformer-based time series forecasting models, placeholders are embedded into a $[1,L_{f},D]$ matrix, $\mathbf{H}_{f}$. The decoder consists of $l$ stacked RACA modules for inputs of length $L_{f}$. Each module has two parallel units: Unit 1, like in Function (1), uses $\mathbf{H}_{f}$ as Query with Key and Value from $\mathbf{X}^{e}_{o}$, outputting $\mathbf{H}^{l,1}_{f}$ of the same shape. Unit 2, retrieval-augment part, like in Function (2) ,also queries $\mathbf{H}_{f}$ but fetches Keys and Values from the concatenated sequence of $\mathbf{X}^{b}_{r}$, generating $\mathbf{H}^{l,2}_{f}$. $s$ in the Functions is the scaling factor.
$\mathbf{H}^{l,1}_{f}=\text{softmax}\left(\frac{\mathbf{H}_{f}\mathbf{X}^{e}_{o}}{s}\right)\mathbf{X}^{e}_{o},$ (1)
$\mathbf{H}^{l,2}_{f}=\text{softmax}\left(\frac{\mathbf{H}_{f}\mathbf{X}^{e}_{r}}{s}\right)\mathbf{X}^{e}_{r},$ (2)
$\mathbf{H}^{l+1}_{f}=\delta\left(\text{concat}\left(\mathbf{H}^{l,1}_{f},\mathbf{H}^{l,2}_{f}\right)\right).$ (3)

### v2

@@
Abstract.
An efficient customer service management system hinges on precise forecasting of service volume. In this scenario, where data non-stationarity is pronounced, successful forecasting heavily relies on identifying and leveraging similar historical data rather than merely summarizing periodic patterns. Existing models based on RNN or Transformer architectures may struggle with this flexible and effective utilization. To tackle this challenge, we initially developed the Time Series Knowledge Base (TSKB) with an advanced indexing system for efficient historical data retrieval. We also developed the Retrieval Augmented Cross-Attention (RACA) module, a variant of the cross-attention mechanism within Transformer’s decoder layers, designed to be seamlessly integrated into the vanilla Transformer architecture to assimilate key historical data segments. The synergy between TSKB and RACA forms the backbone of our Retrieval-Augmented Time Series Forecasting (RATSF) framework. Based on the above two components, RATSF not only significantly enhances performance in the context of Fliggy hotel service volume forecasting but also adapts flexibly to various scenarios and integrates with a multitude of Transformer variants for time-series forecasting. Extensive experimentation has validated the effectiveness and generalizability of this system design across multiple diverse contexts.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
1. Introduction
Accurate estimation of total service demand is crucial for customer service volume management in the travel industry, including hotels, airlines, attractions, and transaction-brokering apps like Fliggy, significantly affecting system costs.Underestimating by 100 service requests incurs urgent mobilization costs, equivalent to three times the labor cost for a single request; overestimation, meanwhile, leads to wasted labor costs. The travel industry presents unique challenges in service volume forecasting due to its interplay with variables like order statuses, weather, international events, and destination country policies. At Fliggy, precise forecasting is essential for planning the recruitment and training of customer service staff, as well as scheduling. These factors result in non-stationary data patterns that can introduce significant biases with traditional uni-variant time-series forecasting methods (Hyndman and Athanasopoulos, 2018; Anderson and Kendall, 1976; Taylor and Letham, 2018), which rely on periodic summaries and trend analyses.
@@ 1. Introduction
One naive way to solve this is to try elongating the sequence that the transformer processes. Some approach involves sampling historical data to fit within a limited context window. Specifically, the Informer (Zhou et al., 2021) algorithm samples K points from the sequence and derives a shorter Q sequence based on these sampled points. However, this approach assumes that all historical information is equally important, which may not be suitable for many time series scenarios where different data points can carry varying significance. Perceiver (Jaegle et al., 2021) and similar methods opt for a different approach by mapping Query sequences to fixed lengths, reducing computation and allowing for more historical data storage. Nonetheless, they still face challenges in efficiently extracting and interpreting critical information from extended time series.
In the field of natural language processing, strides have been made to expand Transformer models’ context handling. On one front, techniques like flash-attention (Dao et al., 2022) enhance efficiency and reduce complexity, enabling longer context processing. However, it’s crucial to note that directly extending receptive fields may cause larger models to overlook significant details in lengthy inputs, like (Liu et al., 2023) mentioned.
On another front, Retrieval-Augmented Generation (RAG)-like methods have drawn attention for enhancing model performance by incorporating external information. NVIDIA’s (Xu et al., 2024) research indicates that even when models already handle large context windows in text tasks, they can still achieve substantial performance gains by retrieving and using relevant data from external sources.
To address the above challenges, we have identified a potential approach employing the concept of retrieval augmentation (RA). Concretely, we focus on implementing two central enhancements: a knowledge base schema that can efficiently index all historical series, and a cross-attention module embedded in a transformer model to integrate historical information for pinpointing and exploiting the most predictive segments, thereby refining prediction accuracy. The synergy between TSKB and RACA constitutes the RATSF framework, which significantly improves the performance of most univariate time series forecasting tasks. Currently, RATSF has been integrated into four key service areas within Fliggy, encompassing hotel bookings and after-sales services, train ticketing, and flight reservations and modifications. This integration has led to a notable decrease in forecasting errors, from an approximate range of 15% to approximately 8% across all the aforementioned sectors, which has significantly reduced personnel management costs, although the exact extent of cost savings cannot be divulged for commercial reasons.
@@
1. We present a straightforward and manageable design for a time-series knowledge base (TSKB) based on the characteristic that time series data is easy to be stored in a structured way, which significantly facilitates efficient management of historical data.
2. We introduce a versatile cross-attention module, retrieval augmented cross-attention (RACA) module, designed to integrate retrieved historical data into the forecasting process. This module is easily adaptable and can be seamlessly integrated with various time-series transformers.
3. We raise a Retrieval-Augmented Time Series Forecasting (RATSF) framework, which decreases 7% forecasting errors in the real service areas within Fliggy and therefore reduced huge personnel practical management costs. Extensive experiments on more publicly available datasets demonstrate that our method achieves the best performance for the univariate time series forecasting task and is general for a broader range of industrial applications.
2. Review
@@
Despite intense competition (Salinas et al., 2020; Das et al., 2023; Ye et al., 2022; Han and Woo, 2022), Transformer models (Vaswani et al., 2017) and their improved variants (Qi et al., 2021; Woo et al., 2022; Liu et al., 2021; Kitaev et al., 2020; Lim et al., 2021) have become the mainstream choice in time-series forecasting tasks. However, the computational complexity of the original Transformer model scales as O(n²) with respect to sequence length, significantly limiting the maximum sequence length it can handle and thereby constraining its ability to use historical data.
@@
Improvements in Time-Series Forecasting with Transformers. Numbers of works focus on enhancing the Transformer model’s capability to extract temporal features, thereby improving prediction accuracy, exemplified by Fedformer’s (Zhou et al., 2022) introduction of a frequency-augmented Attention mechanism that directly incorporates Fourier operators, a similar path taken by TimesNet (Wu et al., 2023). In contrast, Autoformer (Wu et al., 2021) proposes an operator that decomposes time series information into trend and periodic components. These two types of solutions perform well for relatively stationary sequences but may see performance drop when dealing with highly non-stationary time series data.
@@
Another core strategy involves learning a robust representation of the time series first, which is then used to enhance the forecasting accuracy. TNC (Tonekaboni et al., 2021) harnesses the core concept of contrastive learning and employs samples within a specific temporal neighborhood within a window as positive pairs, while treating samples from differing temporal neighborhoods as negative pairs. Despite these improvements having collectively boosted the ability of Transformers in time-series forecasting tasks, none has fundamentally increased the Transformer’s receptive field.
@@
Retrieve Augmented Method in NLP. In last two years, the Retrieval-Augmented Generation (RAG) approach (Guu et al., 2020; Borgeaud et al., 2022; Pasupat et al., 2021; Zhang et al., 2022; Shuster et al., 2021; Khandelwal et al., 2020; Zuo et al., 2019) has gained widespread adoption in the field of NLP. REALM harnesses a knowledge retriever to distill information from vast corpora and thereby enhance the performance of pre-trained language models. Transformer-XL+kNN (Bonetta et al., 2021) incorporates a K-Nearest Neighbors algorithm to search through training data, refining dialogue generation capabilities. In the context of named entity recognition tasks, U-RaNER (Tan et al., 2023) utilizes multi-modal heterogeneous retrieval techniques to boost knowledge retrive and, by integrating retrieved knowledge into the model, strengthens its understanding of queries and improves entity recognition accuracy.
@@
Retrieval Augmented Method in Time-series Forecasting. MQ-ReTCNN (Yang et al., 2022) is designed for complex time series prediction tasks involving multiple entities and variables. It employs a scoring function to compile relevant contexts from offline data, selects scored segments, and appends them to the prediction sequence.Its retrieval mechanism emphasizes leveraging historical sequences of one entity to enhance predictions about another, without directly utilizing historical data for direct prediction assistance.
@@
ReTime (Jing et al., 2022) creates a relation graph based on temporal closeness between sequences and employs relational retrieval instead of content-based retrieval.It does not optimally use historically similar sequences as reference points due to its inherent design limitations. Both MQ-ReTCNN and ReTime incorporate retrieval enhancement strategies but have yet to introduce a general and efficient retrieval technique specifically for single-variable time series prediction scenarios.
Figure 1. RATSF uses a TSKB to store and index historical sequences, alongside a transformer-based forecasting model. As illustrated in the bottom-right corner, TSKB segments the full history for efficient retrieval. To do forecasting, the Encoder fetches $\mathbf{X}_{o}$ from t recent time steps, and forms a retrieval sequence with d latest steps, retrieves N related sequences $\mathbf{X}_{r}$ from the TSKB. RACA in the Decoder then merges $\mathbf{X}_{r}$ and $\mathbf{X}_{o}$ info to deliver result.
3. Method
3.1. Setting & Notations
Before elucidating our approach, we first clarify the setting of our problem and define some symbols that will be used throughout the text.
In actual business operations, service volume is influenced by a multitude of factors, some of which are challenging to fully represent through variables. To simplify the reasoning process and enhance the versatility of our system, we have configured our hotel service volume prediction using a uni-variant time-series forecast setup. This means that the inputs to the model consist solely of the the values of the time series and it’s temporal features . Due to the requirements of our actual task—where personnel management necessitates advance preparation for recruitment, short-term scheduling, and training. Our forecasting task is designed to start from a specific time point $t$ and predict a sequence of $L_{f}$ data points in one go.
Time Marking. We take moment $t$ as the reference point, and the collection of the future time points to be forecasted is represented as ${\left[t+1,t+2,\ldots,t+l_{f}\right]}$, where $l_{f}$ denotes the length of the forecast period. Concurrently, we define the retrieve segment length as $l_{r}$, with the indexing sequence $\mathbf{K}$ in the TSKB having a length of $L_{r}$, and the length of the V sequence as $l_{v}$.
Sequence Source Marking. The subscript o denotes the original sequence, that is, the sequence $\mathbf{X_{o}}={x_{1},...,x_{t}}$ prior to moment $t$; $\mathbf{X_{f}}$ represents the sequence to be forecasted; $\mathbf{X_{r}}$ signifies the $\mathbf{V}$ sequence retrieved from the TSKB.
Transformer Processing Marking. The superscript is used to indicate the results after processing by various parts of the Transformer. $\mathbf{X^{b}_{o}}$ denotes the original sequence after Embedding, and $\mathbf{X^{e}_{o}}$ denotes the sequence after the Encoder. Within the Decoder, H represents the internal hidden state of the Decoder, $l$ represents the layer number, and $H^{l}$ indicates the hidden state processed by the $l$-th layer.
3.2. Overview of RATSF
The RATSF system is composed of two core components: a Time Series Knowledge Base (TSKB, detailed in Section 3.3) and a Transformer-based time series forecasting model. The latter has been enhanced by replacing its original decoder with our novel Retrieval Augmented Cross-Attention (RACA, described in Section 3.4) module. The TSKB efficiently segments and archives historical time series data, establishing a precise indexing system. The transformer model, with the integration of RACA, effectively utilizes retrieved historical data to significantly enhance the accuracy of its predictions. The data flow diagram of RATSF is illustrated in Figure 1, providing a visual representation of the process.
3.3. TSKB
Unlike traditional methods that solely store and retrieve the entire original sequence, TSKB preserves the original sequence as the core content (V) while selects certain segments from it to construct a discriminative indexing sequence (K). This approach can be likened to processing a written piece where conventional methods involve direct full-text searches for desired information, which may be inefficient and less precise; whereas with TSKB, it’s akin to extracting key headlines from the body of the text to serve as indices that facilitate rapid access to core information. This dual-sequence structure allows the system, when handling large-scale data, to efficiently locate and access targeted portions of the original sequence through customized index sequences, thereby significantly enhancing overall performance.
3.3.1. Sequential Slicing
As illustrated in Figure 2, the content sequence V is obtained using a rolling window approach with a step size of $S$ and a window length of $L_{v}$. This approach allows us to sample all historical sequences and incorporate them into TSKB. Since prediction models only use present data to infer the future, we align indexing sequence K with this constraint. We extract the initial segment of each V, having a length of $L_{r}$, to serve as its index K. The selection of $L_{v}$, $L_{r}$ and $S$ is introduced in section 3.5.
Figure 2. TSKB utilizes a rolling window of length $L_{v}$ to collect V, with an indexing segment of length $L_{r}$ taken from its leading part as K , and advances the window in steps of size S.
Figure 3. Within the cold start stage, we use raw sequence of retrieval input to find Top-K relevant sequences using DTW. After one epoch, we switch to Euclidean Distance for retrieval embedding to gather Top-K matches.
3.3.2. Embedding Learning
A key approach to enhancing the recall and accuracy of knowledge base retrieval is to employ a well-trained embedding vector to improve the precision of the index. In the following sections, we will focus on how to train and utilize these index embedding vectors. Judging the quality of a retrieval embedding mainly depends on how well it captures similarities at key information points relevant to forecasting tasks. Time series data is complex due to numerous information points and the challenge of presetting comparative weights. Therefore, we use this principle: if a representation closely mirrors the important details needed for prediction, its overall performance will be better.
To ensure that the embeddings capture the essential information for forecasting, we employ the encoder from the RATSF’s forecasting model, which is designed to select precise information during training, thereby enhancing forecast accuracy. Specifically, each indexing sequence K rom the knowledge base is processed through the encoder layer of the RATSF forecasting model to generate a retrieval embedding.
Meanwhile, in the early stages of training the forecasting model, its encoder is not yet capable of producing representations that accurately match target sequences. The model’s effectiveness and learning pace are significantly correlated with its ability to retrieve historical sequences beneficial to forecasting tasks.
To avoid time and data-consuming iterations stemming from a random initialization state, we introduce Dynamic Time Warping (DTW(Müller, 2007)) as an auxiliary tool during the initial phase of model training, as shown in left part of Figure 3. DTW is initially used for similarity-based sequence retrieval, aiding in the iterative training process of the RATSF forecasting model. After completing one epoch of training, we transition to using embeddings generated by the forecasting model itself for sequence retrieval, continuing the training until the model converges.
@@ 3.3.2. Embedding Learning
Figure 4. Each RACA has two cross-attention modules: one utilizes the Encoder output as K,V, while the other employs embedded retrieved sequences as K,V. Both outputs are concatenated and passed through a Linear module to reshape back to the input dimensions.
3.4. RACA
@@
As previously mentioned, Retrieval-Augmented Cross-Attention (RACA) is a module designed to be integrated into a time-series forecasting model, coupled with the TSKB. In our demonstration, we employ the vanilla Transformer as the time-series forecasting model; however, it’s important to note that the RACA module is designed to be compatible and can be seamlessly plug into any transformer-based time-series model. We will specifically demonstrate this in Section 4 .
@@
The decoder consists of $l$ stacked RACA modules for inputs of length $L_{f}$. Each module has two parallel units: Unit 1, like in Function (1), uses $\mathbf{X}^{b}_{f}$ as Query with Key and Value from $\mathbf{X}^{e}_{o}$, outputting $\mathbf{H}^{l,1}_{f}$ of the same shape. Unit 2, retrieval-augment part, like in Function (2) ,also queries $\mathbf{X}^{b}_{f}$ but fetches Keys and Values from the concatenated sequence of $\mathbf{X}^{e}_{r}$, generating $\mathbf{H}^{l,2}_{f}$. $s$ in the Functions is the scaling factor.
Like shown in Figure 4, retrieved sequences are transformed through the encoder’s embedding module, generating Ns $\mathbf{X}^{e}_{r}$, and we concatenate these N vectors into one with shape$[1,N*L_{v},D]$ for later use.
(1) $\mathbf{H}^{l,1}_{f}=\text{softmax}\left(\frac{\mathbf{X}^{b}_{f}\mathbf{X}^{e}_{o}}{s}\right)\mathbf{X}^{e}_{o},$
(2) $\mathbf{H}^{l,2}_{f}=\text{softmax}\left(\frac{\mathbf{X}^{b}_{f}\mathbf{X}^{e}_{r}}{s}\right)\mathbf{X}^{e}_{r},$
(3) $\mathbf{H}^{l+1}_{f}=\delta\left(\text{concat}\left(\mathbf{H}^{l,1}_{f},\mathbf{H}^{l,2}_{f}\right)\right).$